Which of the following forms the most polar bond?

blank.pngChoose one:

F and C

B and N  

S and Cl

F and Te

Are any of the bonds formed between the following pairs of atoms ionic?

blank.pngChoose one or more:

C and Cl

B and H

Ca and Cl

O and H

P and F
Answer #1

If the elecectronegativity difference b/w the two atoms is <0.4 ; it is pure covalent bond

If the elecectronegativity difference b/w the two atoms is 0.4 to 1.8 ; it is polar  covalent bond

If the elecectronegativity difference b/w the two atoms is > 1.8 ; it is ionic bond

F and C , the EN difference is 3.98- 2.55= 1.43 so it is polar covalent bond

B and N ; the EN difference is 3.04- 2.04 = 1.0 so it is a polar covalent bond

S and Cl; the EN difference is 3.16- 2.58 = 0.58 , it also polar covalent bond

F and Te; 3.98- 2.1 = 1.88 , it not an ionic bond

but and C , the EN difference is 3.98- 2.55= 1.43 so it is most polar covalent bond

Ca and Cl ; EN difference = 3.16- 1.0 = 2.16 ; it is an ionic bond